At low pressure, Vander Waals equation and is reduced to??

Grade:11

1 Answers

Umakant biswal
7 years ago
@ bhavik , 
at low pressure the van der walls equation will be reduced to 
(P+a/VSQUARE )V = RT 
PV+a/V = RT 
PV = RT- a/V  OR PV
this accounts for dip, in PV vs P isotherm at low pressure , 
ALL THE BEST ..